A a series circuit consists of resistor R = 3 XC =5 Omega,XL =9 Omega with a voltage source of E=50V angle O^ and f = 60Hz , Find the reactive power for the capacitor


1
Expert's answer
2021-04-08T08:54:17-0400

"R=3\Omega"

"X_c=5\Omega\\X_L=9\Omega"

Total impedance of the circuit"(Z)" = "\sqrt{R^2+(X_L-X_c)^2}=\sqrt{9+16}=5\Omega"

"V_{RMS}=50V"

"I_{RMS}=\dfrac{V_{RMS}}{Z}=\dfrac{50}{5}=10\space A"

Reactive Power = "I_{RMS}^2Z=100\times5=500\space VAR"